Find the integral of: sec^n (x)dx
1
Expert's answer
2021-01-21T20:07:43-0500

"I_{n}=\int sec^nxdx --------(1)"

"I_{n}=\int sec^{n-2}x.sec^2xdx"

Integration by using parts where as "\int sec^{n-2}xdx" is "I_{st}" and "\int sec^{2}xdx" is "II_{nd}" part


"I_{n}=sec^{n-2}x\int sec^2xdx-\int[d/dx(sec^{n-2}).\int sec^2xdx]dx"

Since "\int sec^2xdx = tanx" and where as "d\/dx(sec^{n-2}x=(n-2)sec^{n-3}xsecxtanx"


"I_{n}=sec^{n-2}x.tanx-\int(n-2)sec^{n-3}x.secx.tanx.tanxdx"

"=sec^{n-2}x.tanx-(n-2)\int sec^{n-2}x.tan^2xdx"

"=sec^{n-2}x.tanx-(n-2)\int sec^{n-2}x(sec^2x-1)dx"

"=sec^{n-2}x.tanx-(n-2)[\int sec^{n}xdx-\int sec^{n-2}xdx"

"=sec^{n-2}x.tanx-(n-2)[\int sec^{n}xdx-\int sec^{n-2}xdx]"

"=sec^{n-2}x.tanx-(n-2)[I_{n}-I_{(n-2)}]"

From equation (1) "I_{n}=\int sec^nxdx" therefore "I_{n-2}=\int sec^{n-2}xdx"


"I_{n}=sec^{n-2}x.tanx-(n-2)I_{n}+(n-2)I_{(n-2)}"

"I_{n}+(n-2)I_{n}=sec^{n-2}x.tanx+(n-2)I_{n-2}"

"(n-1)I_{n}=sec^{n-2}x.tanx+(n-2)I_{n-2}" "I_{n}=[sec^{n-2}x.tanx\/(n-1)]+[(n-2)I_{n-2}\/(n-1)]"

This is the reduction formula of "\int sec^{n-2}xdx"

Answer: "I_{n}=[sec^{n-2}x.tanx\/(n-1)]+[(n-2)I_{n-2}\/(n-1)]"